# Combining straight-line and map-based distances to investigate the connection between proximity to healthy foods and disease

### v1

@@ Abstract
Healthy foods are essential for a healthy life, but accessing healthy food can be more challenging for some people than others. This disparity in food access may lead to disparities in well-being, potentially with disproportionate rates of diseases in communities that face more challenges in accessing healthy food (i.e., low-access communities). Identifying low-access, high-risk communities for targeted interventions is a public health priority, but current methods to quantify food access rely on distance measures that are either computationally simple (like the length of the shortest straight-line route) or accurate (like the length of the shortest map-based driving route), but not both. We propose a multiple imputation approach to combine these distance measures, allowing researchers to harness the computational ease of one with the accuracy of the other. The approach incorporates straight-line distances for all neighborhoods and map-based distances for just a subset, offering comparable estimates to the “gold standard” model using map-based distances for all neighborhoods and improved efficiency over the “complete case” model using map-based distances for just the subset. Through the adoption of a measurement error framework, information from the straight-line distances can be leveraged to compute informative placeholders (i.e., impute) for any neighborhoods without map-based distances. Using simulations and data for the Piedmont Triad region of North Carolina, we quantify and compare the associations between various health outcomes (diabetes and obesity) and neighborhood-level proximity to healthy foods. The imputation procedure also makes it possible to predict the full landscape of food access in an area without requiring map-based measurements for all neighborhoods.
@@ 1 Motivation
Eating healthy foods is critical to childhood development and preventing illnesses in adulthood, including diabetes and obesity (Liu et al., 2000; Harding et al., 2008). However, eating healthy foods is not just a choice; having access to healthy food options also determines what people eat. For some people, predominantly those with disabilities and in low-income or minority communities, healthy foods are not always accessible (Bower et al., 2014; Brucker and Coleman-Jensen, 2017). Access can be hampered by physical factors, like geographic proximity and the availability of public transit (Bennion et al., 2022), and social factors, like structural racism and discrimination (Burke et al., 2018; Odoms-Young and Bruce, 2018). Thus, not just consumption of healthy foods but access to them can impact health (Ahern et al., 2011; Gallegos et al., 2021; Kanchi et al., 2021; Li et al., 2021), and disparities in access can potentially perpetuate disparities in health (Larson et al., 2009).
For public health officials seeking to reduce the burden of diet-related non-communicable diseases, understanding (i) the landscape of food access in a community and (ii) the disease’s connection to food access can be informative. When helping public health officials understand these issues, quantifying food access is an important place to start. To identify low-access, high-risk neighborhoods, food access can be quantified using (i) proximity, i.e., how far away healthy food stores are, or (ii) density, i.e., how many healthy food stores are within some radius. Using geocoded store locations, we focus on neighborhood proximity to healthy foods, defined as the minimum distance between a neighborhood’s “population center” (a population-weighted centroid) and a healthy food store, and the connection between proximity to healthy foods and diagnosed diabetes and obesity prevalence.
@@ 1.1 Drawbacks of Current Distance Calculations
When calculating distances between neighborhoods and healthy food stores, there currently exists a trade-off between (i) computationally simple methods that are less accurate and (ii) more accurate methods that are computationally complex. Euclidean distance, which draws a straight line along the shortest path between two points (the healthy food store location and the neighborhood’s population center) “as the crow flies” is among the former. A recent systematic review of studies in the United Kingdom found that Euclidean distance and circular “buffers” based on it were the most commonly used methods to quantify food access (Titis et al., 2021). (Buffers are commonly circular catchment areas of some specified radius, e.g., one Euclidean mile, from a point of interest.) They are prominent in studies of U.S. food access as well (e.g., Peng et al., 2020; Sánchez et al., 2022). The Haversine formula, which instead draws the shortest possible arc between two points, can offer better accuracy, particularly over long distances, and it is also computationally simple. Still, both of these calculations ignore transportation infrastructure and natural barriers, underestimating proximity to healthy foods and overestimating neighborhood access. These problems could be exacerbated in rural areas with fewer available stores and fewer accessible straight-line routes.
@@
Map-based (or network) distances more accurately quantify neighborhoods’ proximity to healthy foods and more realistically describe the landscape of food access (Bennion et al., 2022), and they can differ substantially from straight-line ones. For the example in Supplemental Figure S1, it can be seen that the estimated driving distance between the historic Reynolda House in Winston-Salem, North Carolina (NC), and a nearby Food Lion grocery store ($2.65$ miles) was $1.6$ times the straight-line distance between them ($1.64$ miles). Further, the straight-line route in Supplemental Figure S1 cuts through a pond, making it impassible for residents of the Reynolda House.
@@
Improved versions of the food access measures like proximity can be calculated from map-based distances (Sharkey et al., 2009; Matsuzaki et al., 2023). The Google Maps API is incredibly powerful for calculations like this, and open-source software in the statistical computing language R (R Core Team, 2023) can integrate the API into a statistical workflow (Kahle and Wickham, 2013; Cooley, 2023). These tools can calculate the map-based travel distances, offering a more accurate snapshot of a neighborhood’s access. Still, making all necessary calculations may not be feasible due to time-intensive computations and monthly limits on free API usage. Even paid options like ArcGIS can quickly become time-intensive, so obtaining map-based distances for entire studies is potentially unrealistic at scale (e.g., to shape policy at the state or national level versus the county).
@@ 1.2 Straight-Line Food Access As a Mismeasured Covariate
With fixed resources, researchers must choose between a small-scale, more accurate study or a large-scale, less accurate one. For example, it would be reasonable to obtain map-based distances for all neighborhoods in one city or county but not in the state. Instead, researchers might sacrifice accuracy to broaden the scope of their study, leading to biased estimates and statistical power loss in the downstream analysis (Carroll et al., 2006). Still, what if accurate, map-based distances can be calculated for some neighborhoods?
The two-phase design (White, 1982), or internal validation study, is commonly used in the measurement error literature (e.g., Giganti et al., 2020; Nab et al., 2021; Lotspeich et al., 2022) and can be applied here. Straight-line distances are collected for all neighborhoods in Phase I. Then, in Phase II, map-based distances are collected for a chosen subset, which can be strategically selected based on any information fully observed in Phase I. For example, Phase II in the Piedmont Triad analysis was chosen via county-stratified random sampling to ensure geographic diversity. Many other sampling designs are possible, including strategic ones to target different goals (e.g., to minimize the variance of the downstream model estimates).
Straight-line and map-based methods seek to measure the same quantity but take different paths to get there. Thus, assuming that straight-line distance underestimates map-based, a measurement error model is a natural way to relate more and less accurate food access variables derived from these distances (like proximity). Let $X$ and $X^{*}$ denote a neighborhood’s proximity to healthy foods based on map-based (error-free) and straight-line (error-prone) distances, respectively. The classical measurement error model (Fuller, 1987) is a popular choice: $X^{*}=X+U$, where the random errors $U$ are assumed to be independent of $X$. Often, $U$ is assumed to follow a mean-zero normal distribution, but when measuring proximity to healthy foods the straight-line measure $X^{*}$ systematically underestimates map-based $X$ by some absolute amount, so $U\leq 0$ necessarily and a different distribution is needed. It may also be desirable to model straight-line $X^{*}$ as systematcally underestimating map-based $X$ by some relative factor instead. A multiplicative error model is also possible: $X^{*}=XW$, where $W$ is once again assumed to be independent of $X$. Multiplicative measurement error models are not as common as additive ones, and there is no most popular distribution to assume for them. The main consideration in selecting a distribution for $W$ is that $0\leq W\leq 1$, since $X^{*}\leq X$. Fortunately, we do not need to specify the measurement error model to use the proposed multiple imputation approach, and the methods work well under either model. In treating the straight-line proximity to healthy foods as an error-prone version of the map-based proximity, a manageable statistical problem replaces an unmanageable computational one.
@@ 1.3 Statistical Modeling With a Mismeasured Covariate
First, suppose that the errors in the straight-line distances were ignored. If the $X^{*}$ were used instead of $X$ to conduct the so-called “naive” analysis of food access and disease prevalence, the model estimates would be biased and inference under-powered (Carroll et al., 2006). Linear regression estimates will be attenuated (Fuller, 1987), while logistic (Barron, 1977) and Poisson regression (Videk and Wong, 1996) estimates can be attenuated or exaggerated. These results echo intuition: Measurement errors must be addressed and analyses corrected to obtain valid inferences about the relationship between food access and health.
@@
Design-based estimators rely primarily on the sampling probabilities describing which neighborhoods were chosen for map-based measurements (i.e., the study design) to overcome the missing data in $X$. Included in this class are the popular inverse probability weighted (IPW) (Horvitz and Thompson, 1952) and augmented IPW/generalized raking estimators (Deville et al., 1993; Robins et al., 1994; Oh et al., 2021a, b; Amorim et al., 2024). Generally, design-based estimators offer better robustness than model-based ones, because they make fewer assumptions about the error mechanism.
Model-based estimators place an extra model on the error mechanism (i.e., the distribution of map-based $X$ given straight-line $X^{*}$). Imputation, maximum likelihood estimation, and regression calibration are popular members of this class (Carroll et al., 2006). Replacing missing $X$ values with predictions based on $X^{*}$ (i.e., imputation) is a promising and popular option because of its ease of implementation and approachability to statisticians and nonstatisticians. So far, multiple imputation has been used to overcome measurement error in binary outcomes (Edwards et al., 2013) or exposures (Cole et al., 2006; Edwards et al., 2015, 2020); in continuous outcomes and covariates (Shepherd et al., 2012); and in a mix of categorical, continuous, or time-to-event variables (Giganti et al., 2020; Han et al., 2021; Pelgrims et al., 2023; Amorim et al., 2024). When assumptions on the error mechanism are correct, model-based estimators can offer better statistical efficiency than design-based ones.
@@ 1.4 Overview
A new statistical approach is proposed to (i) accurately quantify food access on a large geographic scale from incomplete data and (ii) efficiently model its impact on health with less computational strain. This multiple imputation framework is easily generalizable, as adopting other outcome models, adding other fully observed covariates, and incorporating spatial autocorrelation are very straightforward. Also, food access for all neighborhoods in a study can be predicted despite only validating a subset of them, providing a more accurate and comprehensive estimate of the food access landscape. The rest of the paper is organized as follows. In Section 2, we describe the distance calculations and proposed imputation methods. In Sections 3 and 4, the methods are tested in extensive simulations and applied to data on the northwestern region of North Carolina. In Section 5, we discuss our findings and future directions.
@@ 2.1 Model and Data
The neighborhood-level rates of various health outcomes will be modeled in the following way. Specifically, census tracts were adopted as the “neighborhood” unit in for the analysis in Section 4, since that was the smallest geographic scale at which the data were available. Let $Y$ be the model outcome, denoting the number of cases in a neighborhood ($Y\in\{0,1,\dots\}$). This outcome is offset by $Pop$, denoting the neighborhood population ($Pop\in\{Y,Y+1,\dots\})$. The neighborhood’s food access will be measured by $X$, denoting the proximity to healthy foods ($X\in\mathds{R}^{+}$) from the population center, or population-weighted centroid (United States Census Bureau, 2021). “Healthy food stores” are defined in Section 4.2 and can be thought of as grocery stores or other fresh foods sources.
@@
The prevalence ratio for food access, $\exp(\beta_{1})$, is of primary interest. A prevalence ratio of $\exp(\beta_{1})>1$ indicates that worse proximity (i.e., farther distances to the neighborhood’s nearest healthy food store) is associated with higher prevalence of the health outcome, $\exp(\beta_{1})=1$ indicates that proximity had no impact on prevalence of the health outcome, and $\exp(\beta_{1})<1$ indicates that worse proximity led to lower prevalence of the health outcome. $\exp(\beta_{1})>1$.
@@ 2.2 Distance Calculations
Select map-based distances $D_{ij}$ ($i\in\{1,\dots,N\}$, $j\in\{1,\dots,M\}$) are calculated using the mapdist function from the ggmap package in R to query the shortest driving routes between two sets of latitude and longitude coordinates using the Google Maps API (Kahle and Wickham, 2013). In agreement with the API’s Terms of Service, individual distances $D_{ij}$ were not saved; only the derived variable $X_{i}$, computed as the minimum over all $j$, was saved.
@@ 2.3 Overcoming Missing Map-Based Food Access Data
Missing values of map-based proximity, $X$, will be replaced with random draws from their conditional distributions given their error-prone values, $X^{*}$, the log-transformed outcome, $\log(Y)$, and additional fully-observed covariates, ${\mathbf{Z}}$. This conditional distribution is assumed to be normal, with mean $=\alpha_{0}+\alpha_{1}X^{*}+\alpha_{2}\log(Y)+\boldsymbol{\alpha}_{3}^{\top}{\mathbf{Z}}$ and standard deviation $=\sigma$. Estimated parameters for this distribution, denoted by $\hat{\boldsymbol{\alpha}}=(\hat{\alpha}_{0},\hat{\alpha}_{1},\hat{\alpha}_{2},\hat{\boldsymbol{\alpha}}_{3},\hat{\sigma})^{\textrm{T}}$, are obtained by fitting a linear regression “imputation model” – with $X$ as the outcome and $X^{*}$, $\log(Y)$, and ${\mathbf{Z}}$ as the predictors – to the $n$ “queried” neighborhoods without missing data. Note that the analysis model outcome, $\log(Y)$, must be included in the predictor’s imputation model for the imputation model to be congenial (Moons et al., 2006; D’Agostino McGowan et al., 2024). Other ways of including the outcome in the imputation model were considered, but using $\log(Y)$ performed best in simulations (Supplemental Figure S3).
These imputed values are incorporated into a multiple imputation framework to obtain valid standard error estimates for statistical inference about the prevalence ratio of interest, $\exp(\beta_{1})$. Specifically, the following two-step procedure was adopted from prior work (Shepherd et al., 2012) for each iteration $b$ ($b\in\{1,\dots,B\}$).
@@
    Draw imputed values: Impute missing $X$ with random draws $\breve{X}^{(b)}$ from the conditional distribution parameterized by $\hat{\boldsymbol{\alpha}}$. Let $\widetilde{X}^{(b)}$ denote the imputed variable from iteration $b$, where $\widetilde{X}^{(b)}=X$ if map-based measures are non-missing and $\widetilde{X}^{(b)}=$ $\breve{X}^{(b)}$ otherwise.
@@
For the final analysis, the $B$ sets of estimates are pooled according to Rubin’s rules (Rubin, 2004). This imputation approach is implemented in the impPossum function from the possum R package, available on GitHub at https://github.com/sarahlotspeich/possum.

### v2

@@ Abstract
Healthy foods are essential for a healthy life, but accessing healthy food can be more challenging for some people than others. This disparity in food access may lead to disparities in well-being, potentially with disproportionate rates of diseases in communities that face more challenges in accessing healthy food (i.e., low-access communities). Identifying low-access, high-risk communities for targeted interventions is a public health priority, but current methods to quantify food access rely on distance measures that are either computationally simple (like the length of the shortest straight-line route) or accurate (like the length of the shortest map-based driving route), but not both. We propose a multiple imputation approach to combine these distance measures, allowing researchers to harness the computational ease of one with the accuracy of the other. The approach incorporates straight-line distances for all neighborhoods and map-based distances for just a subset, offering comparable estimates to the “gold standard” model using map-based distances for all neighborhoods and improved efficiency over the “complete case” model using map-based distances for just the subset. Through the adoption of a measurement error framework, information from the straight-line distances can be leveraged to compute informative placeholders (i.e., impute) for any neighborhoods without map-based distances. Using simulations and data for the Piedmont Triad region of North Carolina, we quantify and compare the associations between two health outcomes (diabetes and obesity) and neighborhood-level access to healthy foods. The imputation procedure also makes it possible to predict the full landscape of food access in an area without requiring map-based measurements for all neighborhoods.
@@ 1 Motivation
Eating healthy foods is critical to childhood development and preventing illnesses in adulthood, including diabetes and obesity. [1, 2] However, eating healthy foods is not just a choice; having access to healthy food options also determines what people eat. For some people, predominantly those with disabilities and in low-income or minority communities, healthy foods are not always accessible. [3, 4] Access can be hampered by physical factors, like geographic proximity and the availability of public transit, [5] and social factors, like structural racism and discrimination. [6, 7] Thus, not just consumption of healthy foods but access to them can impact health, [8, 9, 10, 11] and disparities in access can potentially perpetuate disparities in health. [12]
The United States Department of Health and Human Services defines five domains of social determinants of health (SDOH): (i) economic stability, (ii) education access and quality, (iii) health care access and quality, (iv) neighborhood and built environment, and (v) social and community context. [13] These domains cover various aspects of peoples’ surroundings (home, work, etc.) that can affect health outcomes. Access to healthy foods falls primarily under domain (iv) since it hinges upon the presence of healthy food retailers, like grocery stores and farmer’s markets, in a person’s environment. Food insecurity is an important, related SDOH, defined as “a household-level economic and social condition of limited or uncertain access to adequate food.” [14] Importantly, a food-insecure household can still have access since the former depends on economic stability (domain i) while the latter is a function of their built environment (domain iv). Here, we focus on food access as an SDOH.
For public health officials seeking to reduce the burden of diet-related non-communicable diseases, understanding (i) the landscape of food access in a community and (ii) the disease’s connection to food access can be informative. When helping public health officials understand these issues, quantifying food access is an important place to start. To identify low-access, high-risk neighborhoods, food access can be quantified using (i) proximity, i.e., how far away healthy food stores are, or (ii) density, i.e., how many healthy food stores are within some radius. Many other food access metrics are also possible, including a binary indicator of at least one healthy food store within some radius, the relative density of healthy to unhealthy food stores within some radius, or the average distance to healthy food stores within some radius. Using geocoded store locations, we focus on neighborhood proximity to healthy foods, defined as the minimum distance between a neighborhood’s “population center” (a population-weighted centroid) and a healthy food store, and the connection between proximity to healthy foods and diagnosed diabetes and obesity prevalence. We chose proximity to healthy foods for its granular ability to quantify access (in miles) rather than as a discretized count or indicator. This choice also avoided the need to specify a buffer zone, and analyses can be sensitive to how these zones are defined (e.g., a $1$-mile buffer versus a $0.5$-mile buffer can lead to different conclusions). [15]
@@ 1.1 Drawbacks of current distance calculations
When calculating distances between neighborhoods and healthy food stores, there currently exists a trade-off between (i) computationally simple methods that are less accurate and (ii) more accurate methods that are computationally complex. Euclidean distance, which draws a straight line along the shortest path between two points (the healthy food store location and the neighborhood’s population center) “as the crow flies” is among the former. A recent systematic review of studies in the United Kingdom found that Euclidean distance and circular buffers based on it were the most commonly used methods to quantify food access. [16] (Buffers are commonly circular catchment areas of some specified radius, e.g., one Euclidean mile, from a point of interest.) They are also prominent in studies of U.S. food access. [17, 18] The Haversine formula, which instead draws the shortest possible arc between two points, can offer better accuracy, particularly over long distances, and it is also computationally simple. Still, Euclidean and Haversine calculations ignore transportation infrastructure (e.g., roads or public bus/train routes) and natural barriers (e.g., rivers or mountains), underestimating proximity to healthy foods and overestimating neighborhood access. These problems could be exacerbated in rural areas with fewer available stores and fewer accessible straight-line routes.
@@
Map-based (or network) distances more accurately quantify neighborhoods’ proximity to healthy foods and more realistically describe the landscape of food access, [5] and they can differ substantially from straight-line ones. For the example in Supplemental Figure S1, it can be seen that the estimated driving distance between the historic Reynolda House in Winston-Salem, North Carolina (NC) and a nearby Food Lion grocery store ($2.65$ miles) was $1.6$ times the straight-line distance between them ($1.64$ miles). Further, the straight-line route in Supplemental Figure S1 cuts through a pond, making it impassible.
@@
Improved versions of the food access measures like proximity can be calculated from map-based distances. [19, 20] The Google Maps API is incredibly powerful for calculations like this, and open-source software in the statistical computing language R can integrate the API into a statistical workflow. [21, 22, 23] These tools can calculate the map-based travel distances, offering a more accurate snapshot of a neighborhood’s access. While we focus on map-based driving distances, map-based distances based on walking or public transportation can also be obtained. These driving distances were calculated from the fastest route under “normal” traffic conditions, although the software also offers further customization (e.g., a 5:00 PM “rush hour” departure time). Still, making all necessary calculations may not be feasible due to time-intensive computations and monthly limits on free API usage. Even paid options like ArcGIS can quickly become time-intensive, so obtaining map-based distances for entire studies is potentially unrealistic at scale (e.g., to shape policy at the state or national level versus the county).
@@ 1.2 Straight-line food access as a mismeasured covariate
With fixed resources, researchers may have to choose between a small-scale, more accurate study or a large-scale, less accurate one. For example, it may be reasonable to obtain map-based distances for all neighborhoods in one city or county but not in the state. Instead, researchers might sacrifice accuracy to broaden the scope of their study, leading to biased estimates and statistical power loss in the downstream analysis. [24] Still, what if accurate, map-based distances can be calculated for some neighborhoods?
The two-phase design, or internal validation study, is commonly used in the measurement error literature and can be applied here. [25, 26, 27, 28] Straight-line distances are collected for all neighborhoods in Phase I. Then, in Phase II, map-based distances are collected for a chosen subset, which can be strategically selected based on any information fully observed in Phase I. For example, Phase II in the Piedmont Triad analysis was chosen via county-stratified random sampling to ensure geographic diversity. Many other sampling designs are possible, including strategic ones to target different goals (e.g., to minimize the variance of the downstream model estimates).
Straight-line and map-based methods seek to measure the same quantity but take different paths to get there. Thus, assuming that straight-line distance underestimates map-based, a measurement error model is a natural way to relate more and less accurate food access variables derived from these distances (like proximity). Let $X$ and $X^{*}$ denote a neighborhood’s proximity to healthy foods based on map-based (error-free) and straight-line (error-prone) distances, respectively. The classical measurement error model is a popular choice: $X^{*}=X+U$, where the random errors $U$ are assumed to be independent of $X$. [29] Often, $U$ is assumed to follow a mean-zero normal distribution, but when measuring proximity to healthy foods, the straight-line measure $X^{*}$ systematically underestimates map-based $X$ by some absolute amount, so $U\leq 0$ necessarily and a different distribution is needed. It may also be desirable to model straight-line $X^{*}$ as systematically underestimating map-based $X$ by some relative factor instead. A multiplicative error model is also possible: $X^{*}=XW$, where $W$ is once again assumed to be independent of $X$. Multiplicative measurement error models are not as common as additive ones, and there is no most popular distribution to assume for them. The main consideration in selecting a distribution for $W$ is that $0\leq W\leq 1$, since $X^{*}\leq X$. Fortunately, we do not need to specify the measurement error model to use the proposed multiple imputation approach, and the methods work well under either model. In treating the straight-line proximity to healthy foods as an error-prone version of the map-based proximity, a manageable statistical problem replaces an unmanageable computational one.
@@ 1.3 Statistical modeling with a mismeasured covariate
First, suppose that the errors in the straight-line distances were ignored. If $X^{*}$ were used instead of $X$ to conduct the so-called “naive” analysis of food access and disease prevalence, the model estimates would be biased and inference underpowered. [24] Linear regression estimates will be attenuated (i.e., biased toward zero), [29] while logistic and Poisson regression estimates can be attenuated or exaggerated. [30, 31] These results echo intuition: Measurement errors must be addressed and analyses corrected to obtain valid inferences about the relationship between food access and health.
@@
Design-based estimators rely primarily on the sampling probabilities describing which neighborhoods were chosen for map-based measurements (i.e., the study design) to overcome the missing data in $X$. This class includes the popular inverse probability weighted (IPW) [32] and augmented IPW/generalized raking estimators. [33, 34, 35, 36, 37] Generally, design-based estimators offer better robustness than model-based ones because they make fewer assumptions about the error mechanism.
Model-based estimators place an extra model on the error mechanism (i.e., the distribution of map-based $X$ given straight-line $X^{*}$). Imputation, maximum likelihood estimation, and regression calibration are popular members of this class. [24] Replacing missing $X$ values with predictions based on $X^{*}$ (i.e., imputation) is a promising and popular option because of its ease of implementation and approachability to statisticians and nonstatisticians. So far, multiple imputation has been used to overcome measurement error in binary outcomes or exposures [38, 39, 40, 41]; in continuous outcomes and covariates [42]; and in a mix of categorical, continuous, or time-to-event variables. [26, 43, 44, 37] When assumptions on the error mechanism are correct, model-based estimators can offer better statistical efficiency than design-based ones.
@@ 1.4 Overview
A new statistical approach is proposed to (i) accurately quantify food access on a large geographic scale from incomplete data and (ii) efficiently model its impact on health with less computational strain. This multiple imputation framework is easily generalizable, as adopting other outcome models, adding other fully observed covariates, and incorporating spatial autocorrelation are very straightforward. Also, food access for all neighborhoods in a study can be predicted despite only validating a subset of them, providing a more accurate and comprehensive estimate of the food access landscape. The rest of the paper is organized as follows. In Section 2, we describe the distance calculations and proposed imputation methods. In Sections 3 and 4, the methods are tested in extensive simulations and applied to data from the Piedmont Triad region in northwestern North Carolina. In Section 5, we discuss our findings and future directions.
@@ 2.1 Model and data
The neighborhood-level rates of various health outcomes will be modeled in the following way. Specifically, census tracts were adopted as the “neighborhood” unit for the analysis in Section 4 since that was the smallest geographic scale at which the data were available. Let $Y$ be the model outcome, denoting the number of cases in a neighborhood ($Y\in\{0,1,\dots\}$). This outcome is offset by $Pop$, denoting the neighborhood population ($Pop\in\{Y,Y+1,\dots\})$. The neighborhood’s food access will be measured by $X$, denoting the proximity to healthy foods ($X\in\mathds{R}^{+}$) from the population center, or population-weighted centroid. [45] “Healthy food stores” are defined in Section 4.2 and can be thought of as grocery stores or other fresh food sources.
@@
The adjusted prevalence ratio for food access, $\exp(\beta_{1})$, is of primary interest. An adjusted prevalence ratio of $\exp(\beta_{1})>1$ indicates that worse proximity (i.e., farther distances to the neighborhood’s nearest healthy food store) is associated with higher prevalence of the health outcome, $\exp(\beta_{1})=1$ indicates that proximity had no impact on prevalence of the health outcome, and $\exp(\beta_{1})<1$ indicates that worse proximity led to lower prevalence of the health outcome.
@@ 2.2 Distance calculations
Select map-based distances $D_{ij}$ ($i\in\{1,\dots,N\}$, $j\in\{1,\dots,M\}$) are calculated using the mapdist function from the ggmap package in R to query the shortest driving routes between two sets of latitude and longitude coordinates using the Google Maps API. [22] In agreement with the API’s Terms of Service, individual distances $D_{ij}$ were not saved; only the derived variable $X_{i}$, computed as the minimum over all $j$, was saved.
@@ 2.3 Overcoming missing map-based food access data
Missing values of map-based proximity, $X$, will be replaced with random draws from their conditional distributions given their error-prone value, $X^{*}$, the log-transformed outcome, $\log(Y)$, and additional fully-observed covariates, ${\mathbf{Z}}$. There can also be extra fully-observed covariates here (e.g., neighborhood gentrification or car ownership) that are more tailored to capturing the error mechanism through the imputation model rather than contributing to the analysis model. This conditional distribution is assumed to be normal, with mean $=\alpha_{0}+\alpha_{1}X^{*}+\alpha_{2}\log(Y)+\boldsymbol{\alpha}_{3}^{\top}{\mathbf{Z}}$ and standard deviation $=\sigma$. Estimated parameters for this distribution, denoted by $\hat{\boldsymbol{\alpha}}=(\hat{\alpha}_{0},\hat{\alpha}_{1},\hat{\alpha}_{2},\hat{\boldsymbol{\alpha}}_{3},\hat{\sigma})^{\textrm{T}}$, are obtained by fitting a linear regression “imputation model” – with $X$ as the outcome and $X^{*}$, $\log(Y)$, and ${\mathbf{Z}}$ as the predictors – to the $n$ “queried” neighborhoods without missing data. Note that the analysis model outcome, $\log(Y)$, must be included in the predictor’s imputation model for the imputation model to be congenial. [48, 49] In Supplemental Section S.1, other ways of including the outcome in the imputation model were considered, but using $\log(Y)$ performed best in simulations (Supplemental Figure S3).
These imputed values are incorporated into a multiple imputation framework to obtain valid standard error estimates for statistical inference about the adjusted prevalence ratio of interest, $\exp(\beta_{1})$. Specifically, the following two-step procedure was adopted from prior work for each iteration $b$ ($b\in\{1,\dots,B\}$). [42]
@@
    Draw imputed values: Impute missing $X$ with random draws $\breve{X}^{(b)}$ from the conditional distribution parameterized by $\hat{\boldsymbol{\alpha}}$. Let $\widetilde{X}^{(b)}$ denote the imputed variable from iteration $b$, where $\widetilde{X}^{(b)}=X$ if the map-based measure is non-missing and $\widetilde{X}^{(b)}=$ $\breve{X}^{(b)}$ otherwise.
@@
For the final analysis, the $B$ sets of estimates are pooled according to Rubin’s rules. [50] This imputation approach is implemented in the impPossum function from the possum R package, available on GitHub at https://github.com/sarahlotspeich/possum.